OpenStudy (anonymous):

The specific heat of gold is 130 C/kgK. How much energy would be required to raise a 200 gram piece from 22 deg C to 30 deg C?

OpenStudy (michele_laino):

here we have to apply the subsequent formula: \[\Large Q = cm\Delta t = 130 \times 0.2 \times \left( {303 - 295} \right)\] since 22 °C degree coreepsonds to 295 K and 30 °C degree corresponds to 303 K

OpenStudy (michele_laino):

200 grams corresponds to 0.2 Kg

OpenStudy (anonymous):

ok! and so we get 208?

OpenStudy (michele_laino):

yes! and the right unit of measure is Cal or Kcal namely we have: 208 C or 208 Kcal

OpenStudy (anonymous):

ohh ok! yay so that is our solution? :O

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay!! thank you!!:D

OpenStudy (michele_laino):

:)